OpenStudy (anonymous):

simplify (p^-8)-3 I got an answer p^24 is this correct

OpenStudy (anonymous):

right one!

OpenStudy (anonymous):

What class are you taking? I do not think this answer is right. I am in Intermediate Algebra, we just did this last week. I do not think you multiply the exponents. I am looking in my book, give me a second to atleast double check to make sure.

OpenStudy (anonymous):

a^-n=1/a^n

OpenStudy (anonymous):

is it \[(p ^{-8})^{-3}\]

OpenStudy (anonymous):

That is the rule for a negative exponent.

OpenStudy (anonymous):

Nickie_S I am in Algebra1B

OpenStudy (anonymous):

if t is as i wrote then we use the law \[(a ^{m})^{n} = a ^{m*n}\] and get \[p ^{(-8)*(-3)} = p ^{24}\]

OpenStudy (anonymous):

thnks Harkirat that is what I go to

OpenStudy (anonymous):

so the question is as I wrote???

OpenStudy (anonymous):

Harkirat that is how I set it up to solve also

OpenStudy (anonymous):

Nickie_S do u hv a different opinion/answer??

OpenStudy (anonymous):

ok jazil then while posting the question type it as follows: simplify (p^-8)^-3 you missed the second carat sign (^)

OpenStudy (anonymous):

I am looking it up, I thought you plus or minus, but the thing is I haven't came across a - and - exponent in my book. I am no expert so I may be wrong.

OpenStudy (anonymous):

thanks Harkirat I sure did miss the second carat sign but u were smart enough to know exactly what I meant

OpenStudy (anonymous):

@ Nickie_s the laws remain same whether indices are positive or negative......

OpenStudy (anonymous):

41. (t^-3)^-6 t^18 I finally found it, I think yeah you multiply the exponents. its right. (See it was a learning question, lol) Sorry.

OpenStudy (anonymous):

see we can even handle it by the law a^-m = 1/a^m ( 1 )^-3 (p^-8)^-3 =(---- ) ( p^8 ) now we will have to use the law when base is fraction \[(p/q)^{-m} = (q/p)^{m}\]

OpenStudy (anonymous):

so it becomes ( p^8 )^3 p^8*3 p^24 (p^-8)^-3 =(---- ) = ------- = ------ = p^24 !!!!!!! ( 1 ) 1^3 1

OpenStudy (anonymous):

So Nickie_S you can see that is a longer and more complicated way of doing it so first method is nice, short, clean... ☺☻☺

OpenStudy (anonymous):

lol, I always go for the long way. I just thought there was something else to it. So its not 1/(p^8)^3 which then turns into p^24. See I thought it was (p^-8)^-3 =p^(-8+-3) =p^-11